OpenStudy (wyattp17):

Graph the circle with center at (-4, 2) and radius 5. Label the center and at least four points on the circle. Write the equation of the circle.

OpenStudy (anonymous):



OpenStudy (anonymous):

What equation do you think you can write about this?

OpenStudy (anonymous):

drawing's a bit off.... there's just one detail I can't shake off haha

OpenStudy (anonymous):

Im gonna add the points in, just wait

OpenStudy (anonymous):

Don't you think that circle's too small? :)

OpenStudy (anonymous):

well I cant make it exact with drawing

OpenStudy (anonymous):

Granted, but the origin should be inside the circle. And it should cross the y-axis, the radius is 5, after all, longer than the distance of the centre from the y axis :D

OpenStudy (anonymous):

Not at all haha

OpenStudy (wyattp17):

whats the equation

OpenStudy (anonymous):

tell us what you've come up with so far @wyattp17

OpenStudy (wyattp17):

im lost bc u guys both are confusing me

OpenStudy (anonymous):

We can't just give you an equation, we need to know that you understand

OpenStudy (anonymous):

Here's how to do it... if you know the centre of the circle, as well as its radius, then the equation will be easy to dish out... say, the center is \(\large (\color{red}h,\color{green}k)\) and the radius is \(\large \color{blue}r\) then the equation of the circle is: \[\Large (x-\color{red}h)^2 + (y-\color{green}k)^2 = \color{blue} r ^2\]

OpenStudy (wyattp17):

then explain

OpenStudy (anonymous):

^

OpenStudy (wyattp17):

ok so its (-4- H)^2 + ( 2-k) 2 = 5

OpenStudy (anonymous):

What? No, lol, don't touch the x's and y's. Your centre should replace (h,k) and not (x,y) And note the square on the r.

OpenStudy (wyattp17):

wait so the -4 and 2 go in the H and K spot ?

OpenStudy (anonymous):

Yup. ^^

OpenStudy (wyattp17):

-4 for H?

OpenStudy (anonymous):

Yes...

OpenStudy (wyattp17):

ok

OpenStudy (wyattp17):

(x--4)^2 + ( y-2)^ 2 = 5 ^2

OpenStudy (anonymous):

--4? What happens when you have two -'s? haha

OpenStudy (wyattp17):

oops typo

OpenStudy (anonymous):

Fix it.

OpenStudy (wyattp17):

(x+4)^2 + ( y-2)^ 2 = 5 ^2

OpenStudy (anonymous):

That's right :) And there's your equation ^^ And there's no way you'd be able to get that into y = _______ form. haha

OpenStudy (wyattp17):

ok is that it ?

OpenStudy (wyattp17):

whats the center

OpenStudy (anonymous):

Well, you could change 5^2 into 25...

OpenStudy (anonymous):

What do you mean what's the centre, it's given in your bloody question haha

OpenStudy (wyattp17):

thx

OpenStudy (anonymous):

Anytime :) Well, not really haha @drumminboy918 wrap it maybe? :)

OpenStudy (wyattp17):

@PeterPan how do i get the points